Let X,, z . 2. Estimate Xn be ind poisson random variobles with parometors P(120 5 Sn = 160) where she Xi, 1-2 ond 1-75
Answer #1

X X2,..., Xog lid. Poisson random variables, with parameter r. Then &def of Xi is given by P (X=k)- & tek and E(X )=4 - V(Xi)hiso A CO - (5562 - P(-10 16 22 5 27/6) - and flac) = 1 e s where 851 and Mo - (x-4)² 26 e - f12) 2.1 25 flasa e >> flz) Bee256 156 22 e dz 20 Set (9) - 52 ( - Tre l o & X 1.96 1.96 2x22 1.96 = 0.7818 Plizos So S160) - 21907 Here, Re-G2 du = eff (u)